OpenStudy (anonymous):

Write 2 equations with 4 and 8 solving it by additions/ subtraction, in 5 steps?

OpenStudy (anonymous):

4+8=12 or 4-8=4

OpenStudy (anonymous):

could you show me how you just came up with that

OpenStudy (anonymous):

it would be 8-4 = 4 not 4-8=4 because then it would -4

OpenStudy (anonymous):

this need to be done in 5 steps I really don't understand math, but you pink have explain it better

OpenStudy (anonymous):

sorry i forgot pinkpanda

OpenStudy (anonymous):

:) we all make mistakes

OpenStudy (anonymous):

I have to write 2 equations between -12 and +12 in 5 steps, I picked 4 and 8 and it have to be done in addition/subtraction. I been trying to figure this out for 3 days and now it is due today. Please show me an example

OpenStudy (anonymous):

1 |you think about the numbers 2| think about equation 3 |addition first subtraction next 4 |4+8=12 5| 8-4=4

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

dont forget to give him a medal